OpenStudy (anonymous):

Can someone help,solve with steps X^3-10x+30

OpenStudy (xxdragoonxx):

\[X^3-10x = X^3-10x+30\]

OpenStudy (xxdragoonxx):

X^3 -10x = Final answer X^3 - 10x + 30

OpenStudy (xxdragoonxx):

Multiply the exponent of X by 3giving \[X^3\]

OpenStudy (xxdragoonxx):

Multiply x and 10 Multiply x and 1. The answer is x

OpenStudy (xxdragoonxx):

10*x evaluates to 10x

OpenStudy (xxdragoonxx):

X^3-10*x evaluates to \[X^3-10*x\] x^3-10x

OpenStudy (xxdragoonxx):

Final answer is \[x^3-10x+30\] x^3-10x+30